OpenStudy (anonymous):

Evaluate (gf)(8) given that f(x)=sq root of x+17 and g(x)=x^2-43 (gf)(8)=?

OpenStudy (anonymous):

g(f(8)) g(sqrt{25}) g(5) 5^2 - 43 25 - 43 g(f(8)) = -18

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem, compositions of functions gave me trouble too.

OpenStudy (anonymous):

is this composition or multiplication?

OpenStudy (across):

\[f(x)=\sqrt{x+17}\]\[g(x)=x^2-43\]\[g(f(x))=(\sqrt{x+17})^2-43=x+17-43=x-26\]\[(8)-26=-18\]

OpenStudy (anonymous):

Composition, f(g(x))

OpenStudy (anonymous):

\[fg(x)\] or \[f\circ g(x)\]?

OpenStudy (anonymous):

oh ok you have the answer then.

OpenStudy (anonymous):

\[f(8)=\sqrt{8+17}=\sqrt{25}=5\] \[g(5)=5^2-43=25-43=-8\]

OpenStudy (anonymous):

25 - 43 = -18, not -8

OpenStudy (anonymous):

One tip, in g(x), replace every "x" term with f(x)